SinA=-4/5 and cosB=7/25 and if none of A,B is in 4th quadrant then find the values of sin(A-B)

Grade:11

3 Answers

Arun
6 years ago
Dear Sudhir
 
sin A = -4/5
cos A = 3/5
 
Cos B = 7/25
Hence
SinB = - 24/25
 
Now
Sin(A -B) = sin A * cos B - cos A sinB
 Put the value and get the answer.
 
 
Regards
Arun (askIITians forum expert)
Akshat Kumar tripathi
6 years ago
We have sin A =-4/5cos B = 7/25 Then, sin B = 24/25cos A = -3/5Now, sin (A-B) = sin A cos B - cos A sin B = -4/5*7/25 + 3/5*24/25 =72/125-28/125 =44/125
B.Sravanthi Reddy
one year ago
If sin alpha = -4/5 & cos beta 7/25 if none of alpha, beta is in 4th quadrant find the values of sin alpha-beta ,cos alpha +beta ,tan alpha +beta